OpenStudy (anisah.21a):

5 x − 7 y = −8 −7 x + 2 y = −20 How to solve for x and y with substitution method

OpenStudy (precal):

5x-7y=-8 isolate x 5x=7y-8

OpenStudy (anisah.21a):

then?

OpenStudy (anisah.21a):

it gives you x=7/5y-8/5

OpenStudy (precal):

\[x=\frac{ 7y-8 }{ 5 }\]

OpenStudy (precal):

yes that is correct

OpenStudy (anisah.21a):

so that is x?

OpenStudy (precal):

now sub in the other equation

OpenStudy (precal):

yes, it is

OpenStudy (anisah.21a):

not right answer

OpenStudy (anisah.21a):

it doesn't give me a long enough response for the x value when i try to submit the answer . ive tried it many times

OpenStudy (precal):

ok are you using math lab?

OpenStudy (precal):

x=(7/5)y-(8/5)

OpenStudy (precal):

try that one

OpenStudy (precal):

are you looking for the solution

OpenStudy (anisah.21a):

yes

OpenStudy (anisah.21a):

not a long response but one number

OpenStudy (precal):

x=4 y=4

OpenStudy (anisah.21a):

how did you get that?

OpenStudy (precal):

There are many ways to do this problem

OpenStudy (precal):

I used my calculator

OpenStudy (anisah.21a):

may i ask, what you did on the calculator?

OpenStudy (anisah.21a):

i'll rate you best response if you explain yourself and not just tell me the answer, thank you.

OpenStudy (precal):

I did explain it. I have to go now